OpenStudy (anonymous):

Using the principle of mathematical induction, prove that \[\sum_{k=1}^{n} k^3 = (\frac{n(n+1)}{2})^2;\ \ \ \ n \geq 1 \]

OpenStudy (sasogeek):

typical question to expect from agd

OpenStudy (perl):

base case is k=1 then show if it is true for n=k , it is true for k+1

OpenStudy (anonymous):

but I'm stuck at the part where you simplify it after the k+1 part to show it is true :(

OpenStudy (anonymous):

I have to reduce the following to something trivial: \[(\frac{(m+1)(m+2)}{2})^2 = (\frac{m(m+1)}{2})^2 + (m+1)^3\]

OpenStudy (mathmate):

RHS: (m(m+1)/2)^2+(m+1)^2 =(m+1)^2[(m/2)^2+(m+1)] =(m+1)^2[m^2+4m+4]/4 =(m+1)^2[(m+2)^2]/4 =[(m+1)(m+2)/2]^2

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

hmm.... in the first step, you have (m+1)^2... is that a typo?

OpenStudy (mathmate):

Yes, it was. [just woke up!]